# Advances in Radiance Field for Dynamic Scene: From Neural Field to Gaussian Field

### v1

@@ 4.6 Discussion
Dynamic scene reconstruction presents significant challenges due to its inherently ambiguous nature, particularly when limited observations are available in monocular settings. These challenges have led to the development of various representation paradigms, each offering different trade-offs between accuracy, efficiency, and flexibility. For rigid objects, the key challenge lies in accurately aligning tracking results from different time stamps into a unified local coordinate system, enabling effective temporal information aggregation. The precision of object localization and tracking therefore becomes fundamental to successful rigid motion reconstruction. While articulated objects consist of rigidly moving parts, their overall motion is non-rigid, making part-level representation more appropriate than object-level approaches. To unify these parts coherently, category-level hierarchical structures are often introduced as priors with specific kinematic constraints. However, developing such kinematic templates is labor-intensive, resulting in templates for only a limited number of articulated motion categories (e.g., humans, quadruped animals). Consequently, template-free reconstruction of articulated motion remains a meaningful yet challenging research direction, with existing approaches showing promise but requiring further quality improvements.
The canonical space approach represents a fundamental design choice where appearance and geometry remain static in a universal reference frame, while neural deformation fields map points between this canonical frame and observed frames. This approach enables effective scene editing with changes propagating through the deformation field, but struggles with sequences exhibiting extreme deformations or topological changes, where maintaining a single coherent canonical space becomes increasingly difficult. Multiple keyframe approaches offer a middle ground by establishing several reference frames rather than relying on a single canonical space. This paradigm relaxes the constraint of finding one universal reference while maintaining temporal coherence across subsets of frames. The extreme case is frame-by-frame optimization in 4D spacetime representation, where each frame functions as its own keyframe. While this approach can achieve high-quality individual frame reconstructions by focusing on per-frame accuracy, it fails to disentangle motion from scene representation, significantly limiting subsequent manipulation capabilities and producing temporally inconsistent results.
Frame-to-frame flow fields bridge adjacent frames through small point displacements, effectively decomposing complex global deformations into more manageable local transformations. This approach eliminates the need for a shared canonical space and handles large deformations by breaking them into incremental steps. However, these local correspondences often lack temporal consistency over long-term sequences as small errors accumulate over time. Point tracking via trajectory fields addresses this limitation by modeling each point’s path as a continuous function of time rather than discrete connections, constraining motion across an object’s entire lifespan and maintaining temporal coherence throughout the video sequence. However, unrestricted trajectory functions may still yield physically implausible motions without proper regularization.
Motion factorization methods decompose complex motion into a limited set of basis trajectories with time-dependent coefficients, effectively capturing shared motion patterns while reducing the solution search space. This transforms the challenging problem of regularizing implicit motion fields into the more intuitive task of regularizing motion coefficients, often leading to more reasonable and physically plausible recovering. In practical applications, hybrid approaches combining multiple representation paradigms yield superior results, particularly for scenes with mixed motion types including rigid objects, articulated entities, and general non-rigid deformations. The learning paradigm also significantly impacts reconstruction quality, with progressive learning strategies—starting from coarse to fine details, incrementally increasing frequency bands, or employing two-stage reconstruction that separates static and dynamic elements—proving effective in handling complex dynamic scenes while avoiding local minima during optimization.
@@ 5 Challenges and Future Trends
Considering the advances discussed, we identify several crucial directions for future research that may have broader impact in this field.
5.1 Efficient and Robust Reconstruction
Neural radiance fields can render photo-realistic views with impressive quality and compact representation, but suffer from high computational costs during training and inference. While 3D Gaussian splatting achieves real-time performance for static scenes with high quality, the modeling of dynamic motion and time-varying geometry remains time-consuming. Developing efficient reconstruction methods for motion, geometry, and appearance simultaneously is a worthy research direction. For casually captured in the wild videos, camera information is typically unavailable, making robust reconstruction under these challenging conditions practically important. Recent advances in directly regressing 3D point maps from 2D image pairs without camera poses have shown promise for static scenes [225, 226], but adaptation to dynamic environments remains underexplored.
5.2 Manipulability and Editability
Manipulation and editing of 2D images and 3D static scenes have seen tremendous progress in recent years. Some methods have attempted to manipulate 4D spacetime representations through scene-level style transfer [227, 228] and object-level operations (removal, addition, repositioning) [107, 106], or by decomposing scenes into static and dynamic regions [229, 230]. However, fine-grained part and pixel-level editing remains challenging. The key to achieving such granularity is accurate point correspondence to propagate edits along the timeline, ensuring temporal consistency. The explicit nature of 3D Gaussian splatting has drawn researchers’ attention to dense tracking [17], showing a potential ability for pixel-level manipulation; however, transforming the unstructured primitives of Gaussian fields into organized hierarchical representations for structured editing of geometry and appearance for general objects still needs significant research attention.
5.3 Scalability
Dynamic scenes can be conceptualized as static environments with time-varying motion fields, yet neither component scales effectively across spatial, temporal, or motion dimensions. Spatially, while neural radiance fields offer compact representation with small memory footprints, they struggle when extended to vast environments like urban-scale scenes. Common solutions employ divide-and-conquer strategies [231, 232, 233], but memory costs increase dramatically with scene size, quickly exceeding available GPU memory. Temporally, current methods typically process sequences spanning only seconds or minutes [234, 38], as extending to hours or days becomes computationally prohibitive due to the linear relationship between scene duration and resource requirements. Motion-wise, achieving stable and accurate long-term dense tracking remains challenging due to occlusions, and dramatic changes in object geometry and appearance over time [235]. Despite recent advances like NeVRF [236] and LVV [237] , reconstructing large-scale long-term dynamic scenes with complex motions still demands substantial research breakthroughs in unified representations that can simultaneously address spatial extent, temporal duration, and motion complexity constraints.
5.4 Reconstrution by Generation
@@
High-quality, comprehensive observations are crucial for dynamic scene reconstruction, yet practical applications typically rely on casually captured monocular data that provides extremely limited visual information, resulting in incomplete reconstructions [75]. While static scene reconstruction has benefited from generative approaches that synthesize invisible or occluded parts using models like LDMs [238], integrating such generative capabilities into 4D dynamic scene reconstruction remains challenging. This integration requires simultaneously maintaining spatial view consistency, temporal coherence, and plausible motion dynamics–a complex multi-constraint problem. Despite recent advances in unconditional, image-based, and text-prompted 4D content generation [239, 240, 241], current methods predominantly output 2D frame sequences without underlying 3D structure [242], thereby failing to provide comprehensive scene representations. The field now faces the critical challenge of effectively conditioning generative models on monocular or sparse multi-view inputs to produce geometrically and temporally consistent 4D volumetric representations that support free-viewpoint rendering and novel trajectory synthesis.
5.5 Large Langage Models
@@
Large Language Models (LLMs) [243] offer powerful semantic priors for dynamic scene understanding [244, 245], complementing visual foundation models through their world knowledge and reasoning capabilities. Despite their potential, integrating LLMs with 4D reconstruction presents significant challenges: high-fidelity reconstruction requires pixel-precise geometry and appearance modeling, while LLMs primarily provide high-level semantic abstractions difficult to align with fine-grained visual features. Recent approaches like Language Embedded 3D Gaussians [246, 247] demonstrate promising directions by incorporating quantized semantic features into explicit scene representations, enabling language-guided editing and querying of 3D content. Future research opportunities lie in developing bidirectional interfaces between LLMs’ symbolic reasoning and the continuous spatiotemporal representations required for dynamic scenes, potentially enabling physics-aware and semantically meaningful scene manipulation and generation.
@@ 6 Conclusion and Outlook
Outlook. While 3D scene reconstruction has achieved remarkable success, the research frontier has shifted toward comprehensive 4D dynamic volume reconstruction–the cornerstone of world models. Powered by advances in neural rendering, generative models, foundation models, and LLMs, we anticipate rapid progress in simultaneously addressing two fundamental challenges: photorealistic geometry and appearance reconstruction, and consistent, physically plausible temporal motion recovering. In summary, 4D dynamic scene reconstruction presents both significant opportunities and challenges, with the ultimate goal of creating high-fidelity digital twins of real dynamic physical environments. In this era of emerging technologies, we hope this survey serves as a valuable foundation to inspire researchers pursuing advances in this transformative domain.
@@ Acknowledgments
This project is partially supported by NSFC (62476075 and 62376080), the Zhejiang Provincial Natural Science Foundation Key Fund of China (LZ23F030003), and the Zhejiang Province Key Laboratory of Optoelectronic Intelligent Imaging and Aerospace Sensing.

### v2

@@ 4.6 Discussion
Dynamic scene reconstruction presents significant challenges due to the complexity of capturing both spatial and temporal variations. To better analyze the diverse approaches in this domain, we propose a unified framework that encapsulates these methods by conceptualizing any dynamic scene as static reference frames with corresponding transformations to target frames. This framework provides a systematic way to categorize reconstruction techniques based on the number of reference frames employed, as illustrated in Fig 5.
For scenes exhibiting rigid or articulated motion, a single canonical space is typically sufficient to represent both geometry and appearance. Similarly, scenes with relatively simple non-rigid deformations can often be reconstructed using a shared canonical space as a common reference. However, as sequences become longer or motions more complex, relying on a single global canonical space becomes increasingly challenging. In such cases, multiple keyframes serving as local references provide a more effective solution, allowing the scene to be reconstructed across several localized spaces. When this local space framework is reduced to just two frames, it transforms into the frame-to-frame flow field reconstruction paradigm, where each reference frame is mapped to its consecutive neighbor using 3D correspondence estimation. Taking this concept further, each frame can serve as its own reference in a frame-by-frame optimization approach, effectively treating the dynamic scene as a continuous 4D spacetime volume. At the finest granularity, per-point tracking establishes reference at the point level, creating individual trajectories for scene elements.
Generally, as the number of reference frames increases and the level of granularity becomes finer, the reconstruction captures more detailed temporal dynamics. However, this enhanced detail comes at the cost of increased computational complexity. This trade-off between fidelity and practical efficiency represents a fundamental consideration when selecting the appropriate reconstruction approach for specific applications.
@@ 5 Challenges and Future Trends
Manipulability and Editability. While substantial progress has been achieved in manipulating and editing 2D images and 3D static scenes, extending these capabilities to 4D spatiotemporal representations remains challenging. Recent approaches have demonstrated promising results in scene-level style transfer [217, 218] and object-level manipulations (removal, addition, repositioning) [95, 94], as well as decomposing complex scenes into static and dynamic components [219, 220]. However, fine-grained part and pixel-level editing in dynamic scenes presents significant difficulties. The critical challenge lies in establishing accurate point correspondence across time to ensure temporal consistency during edit propagation. The explicit representation afforded by 3DGS has recently enabled advances in dense tracking [17], demonstrating potential for pixel-level manipulation; nevertheless, developing structured editing paradigms for inherently unstructured radiance fields remains an open research problem.
Scalability. Dynamic scene reconstruction faces three critical scalability challenges: spatial extent, temporal duration, and motion complexity. Spatially, radiance fields struggle when extended to vast environments like city-level scenes, where memory requirements grow prohibitively with scene size. While divide-and-conquer strategies have been proposed [221, 222, 223], these approaches often struggle with integration and consistency across boundaries, particularly for dynamic scenes. Temporally, computational demands scale linearly with sequence duration, making reconstruction of extended periods (from minutes to days) increasingly prohibitive with current architectures. This challenge is compounded by the difficulties in maintaining robust long-term tracking for motion recovery, as occlusions and dramatic changes in object appearance frequently disrupt correspondence establishment. Despite recent advances, simultaneously addressing spatial scale, temporal extent, and complex non-rigid motion remains an open research challenge requiring fundamental breakthroughs in scene representation and optimization techniques.
@@
Reconstrution by Generation. Dynamic scene reconstruction faces a fundamental challenge: while high-quality results require comprehensive visual data, practical applications often rely on casually captured monocular footage that provides severely limited information, resulting in incomplete reconstructions [68]. Static scene reconstruction has successfully leveraged generative approaches to synthesize invisible or occluded regions using models like Latent Diffusion Models [224], but extending these capabilities to 4D dynamic scenes remains problematic. Such integration demands simultaneous maintenance of spatial view consistency, temporal coherence, and plausible motion dynamics. Despite recent advances in unconditional, image-guided, and text-prompted 4D content generation [225, 226, 227], current methods predominantly produce 2D frame sequences without underlying 3D structure [228], failing to provide comprehensive volumetric representations. The critical research challenge lies in effectively conditioning generative models on partial inputs to produce geometrically accurate and temporally consistent 4D volumes.
@@
Large Language Models. Large Language Models (LLMs) [229] offer powerful semantic priors for dynamic scene understanding [230, 231], complementing visual foundation models through their world knowledge and reasoning capabilities. Despite their potential, integrating LLMs with 4D reconstruction presents significant challenges: high-fidelity reconstruction requires pixel-precise geometry and appearance modeling, while LLMs primarily provide high-level semantic abstractions difficult to align with fine-grained visual features. Recent approaches like Language Embedded 3D Gaussians [232, 233] demonstrate promising directions by incorporating quantized semantic features into explicit scene representations, enabling language-guided editing and querying of 3D content. Future research opportunities lie in developing bidirectional interfaces between LLMs’ symbolic reasoning and the spatiotemporal representations required for dynamic scenes, potentially enabling physics-aware and semantically meaningful scene reconstruction and manipulation.
@@ 6 Conclusion and Outlook
Outlook. While 3D scene reconstruction has achieved remarkable success, the research frontier has shifted toward comprehensive 4D dynamic volume reconstruction. Powered by advances in neural rendering, generative models, foundation models, and LLMs, we anticipate rapid progress in simultaneously addressing two fundamental challenges: photorealistic geometry and appearance reconstruction, and consistent, physically plausible temporal motion recovering. In summary, 4D dynamic scene reconstruction presents both significant opportunities and challenges, with the ultimate goal of creating high-fidelity digital twins of real dynamic physical environments. In this era of emerging technologies, we hope this survey serves as a valuable foundation to inspire researchers pursuing advances in this field.
@@ Acknowledgments
This project is partially supported by NSFC (62476075 and 62376080), the Zhejiang Provincial Natural Science Foundation Key Fund of China (LZ23F030003), and the Zhejiang Key Laboratory of Optoelectronic Intelligent Imaging and Aerospace Sensing.